Jefferson Lab HALL A Logbook for June 07 - All EXCEPT Automatic Entries
Main INDEX
Saturday
207122 06/30/07 23:06 Chiranjib TARGET SCREENS II
207121 06/30/07 23:05 Chiranjib TARGET SCREENS I
207120 06/30/07 22:44 Jack tech - Left Q3 work during polarity change
207119 06/30/07 22:34 atrig Trigger-Download
207118 06/30/07 22:19 folts tech
207117 06/30/07 22:06 A. Camsonne Changed T4 to 2 out of 3
207113 06/30/07 20:49 Boeglin Ozone strip chart
207112 06/30/07 20:48 boeglin Left Arm HV
207111 06/30/07 20:48 boeglin Right Arm HV
207110 06/30/07 20:46 Boeglin Hall A tools screen
207106 06/30/07 19:35 Jack tech - Q3 work earlier in the day
207105 06/30/07 19:29 Boeglin Run 1663: ignore first 300K events
207104 06/30/07 19:19 Boeglin Laft Arm Dipole field drifts
207103 06/30/07 19:12 Chiranjib CODA rebooted
207101 06/30/07 18:53 Chiranjib TARGET SCREENS II
207100 06/30/07 18:52 Chiranjib TARGET SCREENS I
207090 06/30/07 16:08 bogdanw shift summary
207089 06/30/07 16:07 ycwang Target operator
207086 06/30/07 15:57 ycwang HTadiaton/ozone striptchart
207085 06/30/07 15:48 ycwang Radiator temp stripchart
207082 06/30/07 15:14 bogdanw Hall A tools
207081 06/30/07 15:07 ycwang spot++
207079 06/30/07 12:47 Moffit Copper target Operational Restriction
207078 06/30/07 12:16 Moffit Hall A Status
207076 06/30/07 10:18 ycwang High voltage screens
207075 06/30/07 10:15 ycwang HallA Tools
207072 06/30/07 09:32 ycwang Radiator temperature
207069 06/30/07 08:18 ycwang Target screens
207066 06/30/07 08:01 Elena Kh. Shift Summary
207064 06/30/07 07:54 Elena Kh. raster ??
207061 06/30/07 06:27 Elena Kh. spot++ for run 1651
207060 06/30/07 06:22 Elena Kh. Target screens
207058 06/30/07 06:19 Elena Kh. He target
207052 06/30/07 04:58 Elena Kh. spot++ for run 1648
207049 06/30/07 04:27 Elena Kh. spot++ for run 1647
207048 06/30/07 04:15 Elena Kh. Hall A General Tools
207047 06/30/07 04:03 Elena Kh. Radiation chart
207046 06/30/07 04:01 Elena Kh. HRSR HV Screens
207045 06/30/07 04:00 Elena Kh. HRSL HV screen
207044 06/30/07 03:59 Elena Kh. Hall A General Tools and Current monitor
207042 06/30/07 03:49 Elena Kh. Dummy target
207041 06/30/07 03:24 Moffit left Q3 still down.
207040 06/30/07 02:43 Elena Kh. Controlled Access again
207039 06/30/07 02:24 folts tech
207038 06/30/07 01:30 Elena Kh. Controlled Access
207037 06/30/07 01:04 Elena Kh. Target screens
207036 06/30/07 00:40 Elena Kh. Left Q3 went down
207034 06/30/07 00:22 Elena Kh. small left Q3 oscillation
Friday
207031 06/29/07 23:58 Ramesh Shift Summary
207030 06/29/07 23:55 folts tech
207029 06/29/07 23:48 Ramesh Left Q3 back on
207028 06/29/07 23:06 Ramesh Asked for Hall access
207027 06/29/07 22:44 Ramesh left HRS Q3 crashed
207019 06/29/07 20:26 Ramesh right Dipole is up, taking data now
207017 06/29/07 20:19 Ramesh spot++ with carbon (hole) targe
207016 06/29/07 20:18 folts tech
207013 06/29/07 19:32 O. Geagla before run 1640
207012 06/29/07 19:15 Ramesh RHRS dipole still ramping up
207007 06/29/07 16:01 Boeglin Shift Summary
207005 06/29/07 14:49 boeglin Sudden He level drop in Left and Right HRS ESR problem ?
207004 06/29/07 14:43 hyao ESR_FLOW A_4K
207002 06/29/07 14:15 hyao target status
206997 06/29/07 12:35 boeglin Q3 shift at end of run 1635
206994 06/29/07 11:58 R. Michaels re: prescale warnings.
206991 06/29/07 11:06 boeglin ozone strip chart
206990 06/29/07 11:05 boeglin Left arm HV screen
206989 06/29/07 11:03 boeglin Right arm HV screen
206988 06/29/07 11:02 boeglin Hall A Tool screen
206980 06/29/07 07:59 Elena Kh. Shift Summary
206979 06/29/07 07:45 Elena Kh. spot++ for run 1632
206977 06/29/07 05:50 Elena Kh. Target screens
206976 06/29/07 05:28 khem snapshort hv screen
206975 06/29/07 05:26 khem snapshot
206974 06/29/07 05:17 Elena Kh. beam down for at least 1 hour more
206973 06/29/07 04:40 Elena Kh. beam down
206971 06/29/07 04:25 Elena Kh. spot++ for run 1630
206970 06/29/07 04:22 Elena Kh. DT for run 1629 was 7-8%
206967 06/29/07 04:04 Elena Kh. spot++ for run 1629
206966 06/29/07 04:02 Elena Kh. no prescales warnings in the replay and spot++
206964 06/29/07 03:52 Elena Kh. ps factors
206962 06/29/07 03:11 Elena Kh. spot++ for run 1628
206961 06/29/07 03:09 Elena Kh. warnings in replay and spot++
206958 06/29/07 01:59 Elena Kh. spot++ for run 1627
206956 06/29/07 01:50 Elena Kh. Target screens
206955 06/29/07 01:49 Elena Kh. Radiator temperature chart
206954 06/29/07 01:49 Elena Kh. Radiation and ozone charts
206953 06/29/07 01:48 Elena Kh. HRSR screens
206952 06/29/07 01:47 Elena Kh. HRSL screen
206951 06/29/07 01:38 Elena Kh. Rebooted Right HRS / 0 free bytes for iocha12?
206950 06/29/07 01:16 Elena Kh. spot++ for run 1626
206949 06/29/07 01:14 Elena Kh. Hall A General Tools and Current Monitor
206948 06/29/07 01:12 Elena Kh. beam down
206944 06/29/07 00:05 Makis P Shift Summary
Thursday
206940 06/28/07 23:30 Makis P scalers chart
206939 06/28/07 23:30 Makis P left arm HV/right arm HV
206938 06/28/07 23:29 Makis P ozone chart
206937 06/28/07 23:28 Mina Kat spot++ run 1624
206936 06/28/07 23:28 Makis P Beam Current chart
206934 06/28/07 23:20 Ramesh 3He target in
206933 06/28/07 23:18 X. Jiang proceed with data taking. Forget about the hole.
206932 06/28/07 23:15 Mina Kat Radiator in
206930 06/28/07 23:05 Mina Kat Hall A Tools
206928 06/28/07 23:00 Mina Kat spot++ raster size check
206925 06/28/07 22:54 Mina Kat spot++ run 1621
206924 06/28/07 22:53 Ramesh Optics target in
206919 06/28/07 22:22 Mina Kat Beam centered on BPMs A and B
206918 06/28/07 22:03 Makis P MCC wants to do Harps scan
206917 06/28/07 21:58 Makis P Quad tolerance: from +-0.1 to +-0.2%
206916 06/28/07 21:55 Makis P MCC ready to test 50uA in Hall
206915 06/28/07 21:43 Makis P Tuned beam in Hall A
206914 06/28/07 18:32 Makis P MCC Update: 2-6hrs
206913 06/28/07 17:02 Makis P Asked MCC to put Hall in Beam Permit
206912 06/28/07 16:10 atrig Trigger-Download
206911 06/28/07 16:07 Boeglin Shift Summary
206909 06/28/07 14:59 atrig Trigger-Download
206908 06/28/07 14:56 atrig Trigger-Download
206907 06/28/07 13:57 boeglin Left Arm Q3 Q3 power supply fluctuating
206906 06/28/07 11:15 boeglin left arm floor mark camera movement works again
206905 06/28/07 11:12 boeglin Q1R instability, interlock failure message
206900 06/28/07 08:00 Moffit Owl Shift Summary
206895 06/28/07 06:11 Moffit Spectrometers back to 65.32°
206893 06/28/07 05:41 Moffit iocha12 out of free bytes.
206892 06/28/07 05:30 Moffit Checklist snapshots
206891 06/28/07 05:19 ipom Target density Vs. run number
206890 06/28/07 05:19 Moffit Run 1616: spot++
206885 06/28/07 04:48 Moffit Spectrometers moved to 67.32°
206883 06/28/07 04:34 Moffit rtd readback.. cont'd
206882 06/28/07 04:06 Moffit Radiator Temperature is gone.
206879 06/28/07 03:05 Moffit Run 1612: Target Status
206878 06/28/07 03:02 Moffit Run 1612: spot++
206875 06/28/07 02:11 Moffit Beamline HV trip
206873 06/28/07 01:46 Moffit Radiator Out.
206869 06/28/07 00:38 Moffit Ozone increase at 1 pass.
206868 06/28/07 00:21 X. jiang spot++ run 1609
206864 06/28/07 00:08 x. jiang fpp door. just for fun
206863 06/28/07 00:05 Makis Petrat Shift Summary
Wednesday
206860 06/27/07 23:22 Makis P ozone chart and radiator temp
206859 06/27/07 23:19 Makis P Beam current chart and Hall A tools
206858 06/27/07 23:15 Mina Kat spot++ run 1607
206857 06/27/07 23:10 X. jiang changed raster size, to be safe
206854 06/27/07 22:46 Mina Kat spot++, raster magnet current fluctuations
206852 06/27/07 22:34 Makis P raster pattern change for runs 1603-1605
206850 06/27/07 22:27 Mina Kat spot++ run 1605
206849 06/27/07 22:22 Mina Kat spot++ run 1603
206848 06/27/07 22:20 Makis P MCC changed raster size during run 1603!
206841 06/27/07 22:07 Mina Kat spot++ run 1602
206840 06/27/07 22:04 Makis P Run 1602 1st production run for kin6 (E=0.849GeV)
206839 06/27/07 21:56 Makis P Run 1600 is test, 1602 is junk
206837 06/27/07 21:53 Mina Kat spot++ run 1600
206832 06/27/07 21:39 Ramesh 3He target in place
206831 06/27/07 21:36 Mina Kat Radiator in
206827 06/27/07 21:11 Makis P About to start optics run/Raster OFF
206825 06/27/07 21:04 Makis p 50uA beam
206822 06/27/07 20:35 Makis P No usable beam yet
206821 06/27/07 19:48 Ramesh Optics target in place
206820 06/27/07 19:08 Makis P Beam in Hall
206819 06/27/07 18:43 Makis P MCC update
206818 06/27/07 17:45 Makis P HRSs set for new kin/Hall A Tools
206817 06/27/07 17:40 HALL A TOOLS HALL A TOOLS
206816 06/27/07 17:37 Makis P Radiator out
206815 06/27/07 17:36 Mina Kat Triger download
206814 06/27/07 17:35 atrig Trigger-Download
206813 06/27/07 17:26 Makis P Tuned beam in Hall
206812 06/27/07 17:24 Makis P HRS settings for new kinemat
206811 06/27/07 17:00 Makis P Pass change/New HRS settings
206809 06/27/07 16:30 Mina Kat spot++ run 1595
206805 06/27/07 16:05 boeglin Shift Summary
206801 06/27/07 14:28 boeglin Ozone strip chart
206800 06/27/07 14:25 boeglin Radiator strip chart
206799 06/27/07 14:24 boeglin HV Left and Beamline
206797 06/27/07 14:22 boeglin HV Right Arm
206796 06/27/07 14:21 boeglin Hall A tools snapshot
206795 06/27/07 14:14 byung target chart
206794 06/27/07 14:13 byung target is 3He from 1590
206784 06/27/07 10:04 atrig Trigger-Download
206783 06/27/07 10:02 byung target -> optics
206775 06/27/07 08:03 Moffit Owl Shift Summary
206770 06/27/07 06:25 O. Geagla run 1584
206767 06/27/07 06:07 O. Geagla run 1583
206766 06/27/07 06:02 Moffit halog snapshots for checklist
206765 06/27/07 05:56 Moffit Run 1583: Spot++
206762 06/27/07 05:37 Moffit Left Q3 has settled.
206761 06/27/07 05:30 Moffit More info...
206759 06/27/07 05:19 Moffit Oops. Raster off on Helium target.
206758 06/27/07 05:17 eip 1.65 kinematics
206756 06/27/07 05:04 Moffit Radiator In... target to Helium-3
206755 06/27/07 04:55 Moffit Alarm beeps from Left Q3
206753 06/27/07 04:40 Moffit Magnets set for 5D
206752 06/27/07 04:34 O. Geagla run 1580
206751 06/27/07 03:54 atrig Trigger-Download
206750 06/27/07 03:54 Moffit trigger downloaed for 5D
206749 06/27/07 03:51 Moffit Start changing for 5D
206747 06/27/07 03:10 Moffit radiator out... really nice beam
206742 06/27/07 02:27 O. Geagla run 1578
206739 06/27/07 01:34 Moffit run 1577:spot++
206737 06/27/07 01:14 Moffit 6% radiator in, moved to helium target
206735 06/27/07 01:02 Moffit Hall A tools for 5B
206733 06/27/07 00:52 Moffit Spectrometers moved for 5B
206732 06/27/07 00:33 atrig Trigger-Download
206731 06/27/07 00:33 Moffit Trigsetup for 5B
206729 06/27/07 00:18 Moffit Right Q2 p0 beeps
206728 06/27/07 00:14 O. Geagla run 1575
206726 06/27/07 00:11 S. Beck Shift summary
Tuesday
206720 06/26/07 23:26 S. Beck No problem in analysis of 5A !
206719 06/26/07 23:10 Chiranjib TARGET SCREENS II
206718 06/26/07 23:09 Chiranjib TARGET SCREENS I
206717 06/26/07 23:06 Chiranjib OZONE Chart
206716 06/26/07 23:06 Chiranjib Hall A current monitor
206715 06/26/07 23:04 Chiranjib HV
206714 06/26/07 23:03 Chiranjib HV screens
206713 06/26/07 23:02 S. Beck Bad histograms from 5A
206712 06/26/07 23:02 Chiranjib HALL A Con tools/rad chart
206711 06/26/07 23:01 S. Beck Bad histograms from kin 5A
206708 06/26/07 22:59 S. Beck Histograms from new kinematic run look bad
206697 06/26/07 21:40 Chiranjib CODA REBOOTED
206696 06/26/07 20:20 Chiranjib TARGET SCREENS
206695 06/26/07 20:10 Jack Work on Left Q2 and Q3 during the day
206694 06/26/07 18:38 Chiranjib Target ----->> OPTICS
206693 06/26/07 18:35 Chiranjib TARGET IOC reboot
206692 06/26/07 18:34 atrig Trigger-Download
206690 06/26/07 18:09 M. Stevens Tech - ramp doors and smoke removal fans operation restored
206689 06/26/07 17:23 atrig Trigger-Download
206688 06/26/07 16:58 atrig Trigger-Download
206687 06/26/07 16:51 atrig Trigger-Download
206686 06/26/07 16:36 X. jiang TV monitor of target tempreture unit.
206685 06/26/07 16:30 Chiranjib Target status
206684 06/26/07 16:09 atrig Trigger-Download
206683 06/26/07 16:06 folts tech
206682 06/26/07 16:02 boeglin shift summary
206678 06/26/07 12:32 atrig Trigger-Download
206677 06/26/07 09:14 boeglin Hall A going to restricted access
206676 06/26/07 09:13 X. Jiang singles rate before and after target alarm
206672 06/26/07 08:09 cusanno Shift Summary
206669 06/26/07 07:26 cusanno radiator in
206667 06/26/07 06:28 cusanno Radiator out
206664 06/26/07 06:04 cusanno Beamline trip?
206661 06/26/07 04:42 cusanno Radiation Monitor and Ozone
206660 06/26/07 04:41 cusanno Radiator Temperature
206659 06/26/07 04:40 cusanno HV Right Arm
206658 06/26/07 04:40 cusanno HV Left Arm and Beamline
206657 06/26/07 04:37 cusanno Hall A Tools
206655 06/26/07 04:14 garibaldi Carbon target in
206651 06/26/07 03:17 cusanno Beam trip
206647 06/26/07 01:52 cusanno Radiator IN
206645 06/26/07 01:03 cusanno Radiator OUT
206642 06/26/07 00:02 Sparveris Shift summary
Monday
206636 06/25/07 22:16 Chiranjib TARGET SCREENS II
206635 06/25/07 22:15 Chiranjib TARGET SCREENS II
206632 06/25/07 21:23 Chiranjib Target screens after alarm
206631 06/25/07 21:22 Chiranjib Target alarm
206628 06/25/07 20:59 Sparveris spot++
206627 06/25/07 20:52 S. Beck A difference in "R-arm TDCs of L-PMT" histograms
206626 06/25/07 20:39 Sparveris current monitor
206625 06/25/07 20:36 Sparveris radiator temp. screen
206624 06/25/07 20:28 Sparveris radiator in
206621 06/25/07 19:58 Sparveris HV screen
206620 06/25/07 19:50 Sparveris ozone snapshot
206619 06/25/07 19:32 Sparveris Tools screen
206618 06/25/07 19:30 Sparveris production - radiator out
206617 06/25/07 19:30 Chiranjib SPOT ++ ( 1547)
206613 06/25/07 19:14 Chiranjib TARGET ------>>> He-3(LOOP1)
206610 06/25/07 18:53 Chiranjib TARGET ------->>> OPTICS
206609 06/25/07 18:44 Sparveris optics target run
206608 06/25/07 17:14 Sparveris Spectrometers set for new kinematics
206607 06/25/07 17:08 atrig Trigger-Download
206606 06/25/07 16:38 Chiranjib TARGET SCREENS II
206605 06/25/07 16:37 Chiranjib TARGET SCREENS
206604 06/25/07 16:37 Sparveris pass change & controlled access
206602 06/25/07 16:18 Sparveris beam is back
206599 06/25/07 16:02 hyao Day Shift Summary
206595 06/25/07 14:10 hyao ozone screen
206594 06/25/07 14:10 hyao Radiator Temp Screen
206593 06/25/07 14:09 hyao HV Screens
206592 06/25/07 14:07 hyao halla tools screen
206585 06/25/07 11:35 byung target chart
206584 06/25/07 11:34 target he3 tsrget moved to 3He
206583 06/25/07 11:05 hyao spot++ for run 1539
206577 06/25/07 09:36 ipom trigger was downloaded only before run 1536
206576 06/25/07 09:35 ipom trigger was downloaded only before run 1536
206575 06/25/07 09:28 x. jiang triggersetup download,
206574 06/25/07 09:28 atrig Trigger-Download
206572 06/25/07 09:21 hyao HallA Screen for optics run
206567 06/25/07 08:01 cusanno Shift Summary
206564 06/25/07 07:39 cusanno He target in, radiator in
206562 06/25/07 07:23 cusanno optics target in
206561 06/25/07 07:23 cusanno new kinematics settings
206559 06/25/07 06:05 cusanno Radiator out
206555 06/25/07 05:25 cusanno Ozone and Radiator Temperature
206554 06/25/07 05:24 cusanno Radiation Monitor
206553 06/25/07 05:23 cusanno HV Right Arm
206552 06/25/07 05:23 cusanno HV Left Arm and Beamline
206551 06/25/07 05:21 cusanno Hall A Tools
206549 06/25/07 05:01 cusanno radiator in
206547 06/25/07 04:50 O. Geagla run 1530
206546 06/25/07 03:53 cusanno radiator out
206539 06/25/07 01:21 cusanno ps2 changed to 56
Sunday
206531 06/24/07 23:59 S. Beck Shift summary
206530 06/24/07 23:52 ycwang Target screens
206527 06/24/07 23:29 X. Jiang left arm bogie problem again
206524 06/24/07 23:11 atrig Trigger-Download
206520 06/24/07 21:01 S. Beck Trigger histogram for LHRS@105.26
206516 06/24/07 20:45 ycwang HTadiaton/ozone stripchart
206515 06/24/07 20:43 ycwang Radiator temp stripchart
206514 06/24/07 20:39 ycwang High voltage screens
206513 06/24/07 20:38 ycwang High voltage screens
206512 06/24/07 20:35 ycwang Hall A Tools
206508 06/24/07 19:09 A. Camsonne Changed reference run in detectorL and detectorR to run 1514
206505 06/24/07 18:19 S. Beck more histograms from LHRS
206504 06/24/07 18:18 S. Beck differences in histograms of LHRS
206503 06/24/07 17:48 ycwang target screens
206498 06/24/07 16:04 bogdanw shift summary
206492 06/24/07 14:32 K.Allada Target screens
206491 06/24/07 14:26 K.Allada Target moved to 3He - Production resumed.
206487 06/24/07 14:12 bogdanw HRS right HV
206486 06/24/07 14:10 bogdanw HRS left HV
206484 06/24/07 14:00 x, jiang trigersetup output
206482 06/24/07 13:56 atrig Trigger-Download
206481 06/24/07 13:52 Techs Replaced middle bogie motor #4
206480 06/24/07 13:51 bogdanw kinematics 12B
206479 06/24/07 11:33 X. Jiang Shift leaders please read
206478 06/24/07 11:06 bogdanw time for ozone to drop
206477 06/24/07 08:41 K.Allada Target Moved to empty as per MCC request
206474 06/24/07 08:03 cusanno Shift Summary
206468 06/24/07 06:24 cusanno Radiator OUT
206466 06/24/07 05:49 E. Cisbani cryotarget charts
206465 06/24/07 05:44 E. Cisbani Target controls screens
206460 06/24/07 04:24 cusanno Radiation Monitors and Ozone
206459 06/24/07 04:22 cusanno Radiator Temperature
206458 06/24/07 04:22 cusanno HV Left Arm
206457 06/24/07 04:20 cusanno HV Right Arm
206456 06/24/07 04:19 cusanno Hall A Tools
206455 06/24/07 03:37 cusanno CODA problem
206450 06/24/07 01:30 cusanno radiator in
206448 06/24/07 00:47 cusanno HRS Left Q3 small drifts
206446 06/24/07 00:08 X. Jiang started C12 run. cm=90 deg.
206445 06/24/07 00:01 Sharon Beck Shift summary
206443 06/24/07 00:01 byung target C
Saturday
206442 06/23/07 23:56 S. Beck Total events He3@2.5 GeV
206437 06/23/07 22:44 S. Beck Raster histograms
206430 06/23/07 19:46 S. Beck He level in left spect. too low
206429 06/23/07 19:31 byung target chart
206424 06/23/07 18:04 S. Beck A correction to 206422 (good run 1496, bad runs 1493,1494)
206423 06/23/07 17:53 X. Jiang Radiator tempreture screen shot.
206422 06/23/07 17:44 S. Beck run 1496: too high temperature on radiator - JUNK
206421 06/23/07 17:40 X. Jiang adjustment on radiator position.
206420 06/23/07 17:39 X. Jiang run 1494 is a junk run.
206417 06/23/07 17:24 X. jiang Hall A tool screen shot.
206416 06/23/07 17:22 X. Jiang spot++ check for run 1494
206413 06/23/07 16:56 byung target chart - HPH
206408 06/23/07 16:01 bogdanw shift summary
206407 06/23/07 16:01 bogdanw radiator at 45 mkA
206403 06/23/07 15:13 gilman 45 uA on radiator
206396 06/23/07 13:36 K.Allada Run 1487 VDC plots
206389 06/23/07 10:45 bogdanw HV left nad T radiator
206388 06/23/07 10:44 bogdanw HV Right
206387 06/23/07 10:44 bogdanw radiation and ozon
206386 06/23/07 10:43 bogdanw Hall A tools and beam
206385 06/23/07 09:57 bogdanw raster for run 1484
206382 06/23/07 08:56 bogdanw radiator out
206378 06/23/07 08:10 cusanno Co2 pressure level low and going down
206377 06/23/07 08:07 cusanno Shift Summary
206374 06/23/07 07:34 cusanno Beam is back
206372 06/23/07 07:25 cusanno beam trip
206370 06/23/07 07:01 E. Cisbani cryotarget charts
206369 06/23/07 07:01 E. Cisbani cryotarget controls screenshot
206366 06/23/07 05:42 A. Camsonne Odd behavior of modules since reboot
206365 06/23/07 05:26 cusanno RMonitor and Ozone
206364 06/23/07 05:25 cusanno Radiator stripchart
206363 06/23/07 05:24 cusanno HV Right Arm
206362 06/23/07 05:23 cusanno HV left arm
206361 06/23/07 05:22 cusanno Hall A Tools
206360 06/23/07 05:16 cusanno radiator IN
206358 06/23/07 05:11 cusanno FPP Threshold found at zero
206356 06/23/07 04:50 A. Camsonne Left Arm DAQ VDC problem
206355 06/23/07 04:36 mcanan VDC problem fixed
206352 06/23/07 04:21 cusanno Beam permit
206351 06/23/07 04:13 cusanno Left VDC fixed
206343 06/23/07 03:28 cusanno Controlled Access
206337 06/23/07 03:04 mcanan VDC eff. run 1464
206336 06/23/07 03:02 mcanan VDC eff. earlier run 1462
206335 06/23/07 02:41 X. Jiang VDC problem cont.
206334 06/23/07 02:38 cusanno Left VDC
206333 06/23/07 02:36 Jack VDC Gas System
206323 06/23/07 01:37 cusanno Left VDC running out of Argon
206319 06/23/07 01:03 mcanan LEFT VDC U1 and U2 wiremaps empty
206318 06/23/07 01:01 cusanno CODA and Left VDC
206316 06/23/07 00:44 cusanno NMR Left Arm
206315 06/23/07 00:43 cusanno VDC bad on replay again
206314 06/23/07 00:33 cusanno Radiator OUT
206312 06/23/07 00:20 cusanno Left VDC is bad on replay
206311 06/23/07 00:01 X. Zheng swing shift summary
Friday
206308 06/22/07 23:43 X. Zheng LeCroy HV IOC reboot and VDC HV reset
206304 06/22/07 22:32 Chiranjib TARGET SCREENS II
206303 06/22/07 22:31 Chiranjib TARGET SCREENS I
206302 06/22/07 21:47 Sparveris junk runs: 1460 & 1461
206298 06/22/07 20:43 X. Zheng radiator IN
206296 06/22/07 19:58 X. Zheng ozone and radiation monitor stripcharts
206295 06/22/07 19:57 X. Zheng radiator temperature stripchart
206294 06/22/07 19:56 X. Zheng Left Arm HV screen
206293 06/22/07 19:56 X. Zheng Right Arm HV screens
206292 06/22/07 19:55 X. Zheng Tools screen and BCM chart
206291 06/22/07 19:53 X. Zheng spot++ for run 1458
206289 06/22/07 19:43 X. Zheng radiator OUT
206283 06/22/07 17:35 X. Zheng hole (3D)
206282 06/22/07 17:11 X. Zheng hole from run 1452
206279 06/22/07 16:35 X. Zheng wrong entry time?
206278 06/22/07 16:25 Jack Work done during the access today
206277 06/22/07 16:20 Chiranjib TARGET SCREENS II
206276 06/22/07 16:18 Chiranjib TARGET SCREENS I
206275 06/22/07 16:16 X. Zheng radiator in since 15:47 (run 1454)
206274 06/22/07 16:11 cusanno HRS-Right Q1 large oscillation
206271 06/22/07 16:01 Yi Qiang Day shift summary
206269 06/22/07 15:49 Yi Qiang Radiator OUT
206267 06/22/07 15:40 Spot++ for r Spot++ 1452
206265 06/22/07 14:41 hyao target=he3
206263 06/22/07 14:19 Xiaohui Spot++
206261 06/22/07 14:13 hyao target=carbon hole
206259 06/22/07 14:07 hyao target chart
206257 06/22/07 13:40 hyao target=optics
206254 06/22/07 13:21 Xiaohui Spot++
206241 06/22/07 12:49 hyao target=carbon
206234 06/22/07 12:14 Yi Qiang Kinematics changed to A12
206233 06/22/07 10:16 A. Camsonne Power cycled TS0
206232 06/22/07 10:14 atrig Trigger-Download
206231 06/22/07 10:14 musson Downstream BCM
206230 06/22/07 08:42 hyao Target=Empty
206228 06/22/07 08:10 gilman radiator temperature during stable beam
206225 06/22/07 08:01 E. Cisbani Owl shift summary
206223 06/22/07 07:13 idar target charts
206222 06/22/07 07:12 idar target charts
206220 06/22/07 07:01 E. Cisbani Radiator OUT
206214 06/22/07 04:51 E. Cisbani Downstream BCM drift recovered
206213 06/22/07 04:27 E. Cisbani Radiation Monitor and Ozone
206212 06/22/07 04:26 E. Cisbani High Voltage Screens
206211 06/22/07 04:25 E. Cisbani Radiator Temperature
206210 06/22/07 04:24 E. Cisbani Hall A General Tools and current monitor
206209 06/22/07 04:00 E. Cisbani Radiator IN (pos. 5)
206204 06/22/07 02:02 E. Cisbani Radiator OUT
206199 06/22/07 00:35 E. Cisbani LEFT and RIGHT momenta differ by 1e-5
206197 06/22/07 00:00 Garibaldi Hall A tools swing shift
Thursday
206196 06/21/07 23:59 Garibaldi Shift Summary
206193 06/21/07 23:42 O. Geagla Run 1431
206188 06/21/07 21:41 Garibaldi Beam trips
206186 06/21/07 20:39 X. Jiang controlled access to fix Q1 problem
206185 06/21/07 19:30 X. jiang right arm Q1 can not reset.
206184 06/21/07 19:12 X. Jiang Right arm Q1 not stable.
206179 06/21/07 17:34 F. Garibaldi Q1 unstable
206173 06/21/07 16:01 E. Schulte Day shift Log
206171 06/21/07 15:15 romanip Target Screens, Vacuum Chart
206168 06/21/07 15:04 eschulte HRs-R Q1 Drifting again
206165 06/21/07 14:52 eschulte HRS-R Q1 drift: Run 1422
206164 06/21/07 14:03 romanip Target Screens
206158 06/21/07 12:25 eschulte Day Shift Checklist: Radiation Monitors & Ozone
206157 06/21/07 12:21 eschulte Day Shift Checklist: HRS-L and HRS-R HV Screens
206156 06/21/07 12:18 eschulte Day Shift Checklist: Hall A Tools
206155 06/21/07 12:13 eschulte Day Shift Checklist: Radiator Temp.
206153 06/21/07 12:02 eschulte Changed to radiator in
206151 06/21/07 11:52 romanip Target Screens, radiator - OUT
206150 06/21/07 11:17 romanip Target Screen, Vacuum Chart
206145 06/21/07 10:03 romanip Target Screens, radiator - OUT
206144 06/21/07 09:58 eschulte BCM: Expert coming out
206139 06/21/07 08:12 romanip Target Screens
206138 06/21/07 08:05 E. Cisbani Owl shift summary
206134 06/21/07 07:31 E. Cisbani 3rd significant drift of Downstream BCM
206131 06/21/07 05:52 E. Cisbani Radiator IN
206128 06/21/07 05:43 E. Cisbani Down - Up stream BCM discrepancy
206125 06/21/07 04:38 E. Cisbani Radiator temperature
206124 06/21/07 04:37 E. Cisbani ozone and radiation monitor
206123 06/21/07 04:36 E. Cisbani HV monitor
206122 06/21/07 04:35 E. Cisbani Hall A Tools and beam stripchart
206121 06/21/07 04:26 Xinhu Yan Target Screens
206119 06/21/07 04:08 E. Cisbani Radiator is OUT
206115 06/21/07 02:34 E. Cisbani Downstream BCM recovering
206112 06/21/07 02:19 E. Cisbani Oscillation (drift) in beam current monitor
206111 06/21/07 01:37 E. Cisbani Continue to run with radiator IN
206105 06/21/07 00:11 Xinhu Yan Target Snapshot
206104 06/21/07 00:09 hyao Swing Shift Summary
Wednesday
206102 06/20/07 23:45 ycwang Target screens
206099 06/20/07 23:20 hyao ozone stripchart
206098 06/20/07 23:17 hyao radiator temp screen
206097 06/20/07 23:16 hyao Magnet Screen
206096 06/20/07 23:12 hyao Hall A tools
206093 06/20/07 21:29 ycwang Target screens
206092 06/20/07 21:28 hyao radiator in
206090 06/20/07 21:25 hyao Radiator=Out
206088 06/20/07 20:35 hyao Radiator Out and Beam Current 48uA
206086 06/20/07 20:22 x. jiang target screen shot
206083 06/20/07 20:17 hyao spot++ for run 1400
206081 06/20/07 19:55 hyao CODA froze
206079 06/20/07 19:33 ycwang Target screens
206078 06/20/07 19:03 ycwang Target screens
206071 06/20/07 16:38 ycwang Target screens
206069 06/20/07 16:00 E. Schulte Day Shift Log
206068 06/20/07 15:34 R. Michaels target snapshot
206066 06/20/07 15:31 eschulte Changed to radiator in
206064 06/20/07 14:40 R. Michaels CODA crash/recovery
206062 06/20/07 14:38 eschulte DAQ: TS0 died on run 1394
206061 06/20/07 14:04 eschulte Day Shift Checklist: Radiation Monitors & Ozone
206059 06/20/07 13:43 eschulte MCC: Back in business.
206058 06/20/07 13:06 eschulte MCC: Spectrometer magnet has a water leak
206057 06/20/07 13:01 eschulte MCC: Spectrometer tripped
206056 06/20/07 12:57 eschulte Inadvertant HRS-L Dipole excersion
206054 06/20/07 12:54 eschulte Changed to radiator out
206052 06/20/07 12:42 eschulte Day Shift Checklist: Radiator Temp.
206051 06/20/07 12:39 eschulte Day Shift Checklist: High Voltage Screens
206050 06/20/07 12:38 eschulte Day Shift Checklist: Hall A Tools
206048 06/20/07 11:53 eschulte eschulte
206047 06/20/07 11:49 eschulte RHS-R Q2 dirft again.
206043 06/20/07 11:43 eschulte HRS-R Q2 drift.
206042 06/20/07 10:54 eschulte latches cleared without incident
206040 06/20/07 10:50 eschulte Paged J. Segal
206035 06/20/07 08:43 eschulte Back to 3He running only.
206029 06/20/07 08:01 E. Cisbani Owl shift summary
206024 06/20/07 05:06 E. Cisbani radiation monitor and ozone stripchart
206023 06/20/07 05:05 E. Cisbani Radiator Temperature stripchart
206022 06/20/07 05:04 E. Cisbani High Voltage Screens
206021 06/20/07 05:02 E. Cisbani Hall A tools
206020 06/20/07 05:00 E. Cisbani TS0 frozen, reboted
206008 06/20/07 00:08 Xinhu Yan Target Screen
206006 06/20/07 00:01 cusanno Shift Summary
Tuesday
206003 06/19/07 23:28 cusanno HRS-Right Q1 large oscillation
206000 06/19/07 21:27 cusanno Carbon with radiator IN
205998 06/19/07 21:16 cusanno Carbon target IN
205996 06/19/07 20:31 cusanno Beam trip
205995 06/19/07 20:21 cusanno ozone and RM strips
205994 06/19/07 20:19 cusanno HV screen 2
205993 06/19/07 20:19 cusanno HV screen 1
205992 06/19/07 20:18 cusanno Hall A tools
205989 06/19/07 19:45 cusanno target screen 2
205988 06/19/07 19:44 cusanno Target screen 1
205987 06/19/07 19:40 A. Camsonne Missing runcontrol
205986 06/19/07 19:30 cusanno Left Q3
205983 06/19/07 18:58 cusanno CODA hiding fixed
205982 06/19/07 18:44 cusanno CODA mistake
205979 06/19/07 17:45 cusanno temporary minor problem with Right Q2
205975 06/19/07 16:03 E. Schulte Day Shift Log
205971 06/19/07 14:53 eschulte Changed to radiator out
205965 06/19/07 12:51 eschulte Day Shift Checklist: Ozone & Rad. Monitors
205964 06/19/07 12:50 eschulte Day Shift Checklist: Radiator Temp
205963 06/19/07 12:49 eschulte Day Shift Checklist: High Voltage HRS-L
205962 06/19/07 12:49 eschulte Day Shift Checklist: High Voltage HRS-R
205961 06/19/07 12:48 eschulte Day Shift Checklist: Hall A Tools
205960 06/19/07 12:34 eschulte Target Screens 1
205959 06/19/07 12:13 eschulte HRS-L VDC v1 wires
205955 06/19/07 11:44 eschulte Beam back, taking data
205944 06/19/07 10:31 A. Camsonne rebooting adaql2
205940 06/19/07 10:09 eschulte MCC: Sending CW
205939 06/19/07 10:07 A. Camsonne Dead time issue : not seen on cosmics so far
205934 06/19/07 09:53 eschulte Moved target and radiator out
205930 06/19/07 08:53 rshneor target density drop effect on singles rates
205925 06/19/07 08:02 X. Zheng owl shift summary
205922 06/19/07 06:34 O. Geagla right arm s2 tdc (again)
205919 06/19/07 06:04 X. Zheng radiator in
205914 06/19/07 04:52 O. Geagla target screen
205911 06/19/07 04:36 X. Zheng Right arm HV screens
205910 06/19/07 04:36 X. Zheng Left Arm HV screen
205909 06/19/07 04:34 X. Zheng Radiator temperature stripchart and BCM monitor
205908 06/19/07 04:31 X. Zheng radiation monitor stripchart and ozone
205907 06/19/07 04:30 X. Zheng Hall A Tools Screen
205902 06/19/07 01:52 X. Zheng spot++ for run 1347
205899 06/19/07 01:06 O. Geagla left arm s1 tdc possible abnormality
205898 06/19/07 01:02 O. Geagla right arm s2 tdc abnormal (maybe)
205895 06/19/07 00:07 A. Camsonne CODA trouble to be studied
205893 06/19/07 00:00 cusanno Shift Summary
Monday
205890 06/18/07 23:18 cusanno CODA problems
205885 06/18/07 22:23 ycwang Target screens
205884 06/18/07 21:35 Adaq target density drop
205879 06/18/07 20:13 ycwang Target screens
205876 06/18/07 19:47 cusanno Ozone strip chart
205875 06/18/07 19:46 cusanno HRS-R HV
205874 06/18/07 19:46 cusanno HRS-L HV
205873 06/18/07 19:44 cusanno RM and Radiator Temp
205872 06/18/07 19:42 cusanno Hall A Tools
205867 06/18/07 18:23 cusanno CODA minor trouble
205865 06/18/07 18:17 ycwang Target screens
205862 06/18/07 17:18 cusanno Radiator off
205860 06/18/07 16:27 cusanno beam tuning in a while
205859 06/18/07 16:12 ycwang Target screens
205856 06/18/07 16:01 eschulte Day Shift Log
205854 06/18/07 15:46 romanip Target Screens
205851 06/18/07 14:19 eschulte Hall A Tools after Q1 Cryo Trip
205849 06/18/07 14:15 eschulte HRS-R Q1 reset
205848 06/18/07 14:11 eschulte Back to Beam Permit
205847 06/18/07 14:11 sawatzky Left Q3 drifting is a known problem
205846 06/18/07 14:09 eschulte HRS-R Q1 resetting
205845 06/18/07 13:41 eschulte HRS-R Q1 needs to be reset in the Hall
205844 06/18/07 13:40 romanip Target Screens
205843 06/18/07 13:33 eschulte HRS-R Q1 Cryo fault
205841 06/18/07 13:24 eschulte HRS-L Q3 Drifted again
205838 06/18/07 13:16 eschulte HRS-L Q3 Drifting low
205837 06/18/07 13:08 eschulte Unidentified UPS-sounding Alarm
205835 06/18/07 12:22 eschulte Changed to radiator in
205833 06/18/07 11:58 eschulte BPM's at Mid Day Shift
205832 06/18/07 11:58 eschulte Day Shift Checklist: RM and Ozone
205831 06/18/07 11:57 eschulte Day Shift Checklist: Radiator Temp
205830 06/18/07 11:56 eschulte Day Shift Checklist: HRS-L HV Screen
205829 06/18/07 11:56 eschulte Day Shift Checklist: HRS-R HV Screens
205828 06/18/07 11:55 eschulte Day Shift Checklist: Hall A Tools
205827 06/18/07 11:50 romanip Target Screens
205825 06/18/07 11:18 eschulte Changed to radiator out
205823 06/18/07 10:41 eschulte Incorrect label for runs 1320 through 1329
205822 06/18/07 10:22 romanip alarm LOW settings for haHe3_P_Fill_Target_R
205820 06/18/07 10:08 romanip Target Screens
205819 06/18/07 10:08 eschulte Hall back to beam permit
205818 06/18/07 10:06 eschulte Target: Pressure drift real
205817 06/18/07 09:31 eschulte Hall to go to Controlled Access
205815 06/18/07 09:21 eschulte MCC: Adjusted Box Supply
205814 06/18/07 08:48 eschulte Target: Controlled Access Scheduled
205811 06/18/07 08:16 romanip Target Screens
205809 06/18/07 07:55 bubis Owl Shift
205808 06/18/07 07:39 Elena Kh. LOW haHe3_P_Fill_Target_R -> 158.00
205805 06/18/07 07:21 bubis ozone
205804 06/18/07 07:19 bubis Radiator temp.
205803 06/18/07 07:18 bubis HV
205802 06/18/07 07:17 Elena Kh. LOW for haHe3_P_Fill_Target_R from 160 to 159
205801 06/18/07 07:17 bubis Hall A Tools
205798 06/18/07 06:01 Elena Kh. Target Screens
205793 06/18/07 04:33 Elena Kh. Low alarm for haHe3_P_Fill_Target_R
205788 06/18/07 02:31 Elena Kh. Target screens
205782 06/18/07 00:00 E. Cisbani Swing shift summary
Sunday
205779 06/17/07 22:54 A. Camsonne et_start cleaned up
205774 06/17/07 20:50 A. Camsonne Target screens
205773 06/17/07 20:48 A. Camsonne Target screens
205770 06/17/07 20:40 E. Cisbani High Voltages snaposhot
205769 06/17/07 20:39 E. Cisbani Radiation and Ozone Stripcharts
205768 06/17/07 20:37 E. Cisbani Radiator Temperature stripchart
205767 06/17/07 20:37 E. Cisbani Hall A Tools Snapshot
205766 06/17/07 20:27 A. Camsonne moved file from /root to /adaql10/work1/rootadaql2
205765 06/17/07 19:47 A. Camsonne Run 1314 failed to start, / partition full on adaql2
205754 06/17/07 16:07 E. Schulte Day Shift Log
205751 06/17/07 14:53 K. Allada Target screens
205749 06/17/07 14:05 eschulte Radiator in again
205747 06/17/07 13:51 gilman FPP thresholds
205746 06/17/07 13:03 eschulte BPM & Current Monitors at Mid-shift
205745 06/17/07 13:01 eschulte Ozone and Radiation: Day Shift Checklist
205744 06/17/07 13:00 eschulte Radiator Temp: Day Shift Checklist
205743 06/17/07 12:59 eschulte HV: HRS-R
205742 06/17/07 12:58 eschulte HV: HRS-L
205740 06/17/07 12:58 eschulte Day Shift Checklist: Hall A Tools
205732 06/17/07 09:00 eschulte Radiator In
205729 06/17/07 08:05 A. Camsonne Owl Shift Log
205725 06/17/07 06:51 A. Camsonne Beam position and current : good beam
205720 06/17/07 05:17 A. Camsonne Radiator temperature
205719 06/17/07 05:16 A. Camsonne Radiation chart / Ozone
205718 06/17/07 05:15 A. Camsonne HV Right arm
205717 06/17/07 05:14 A. Camsonne HV Left arm
205716 06/17/07 05:13 A. Camsonne Hall A tools
205713 06/17/07 03:53 A. Camsonne Radiator temperature
205711 06/17/07 03:43 A. Camsonne Radiator in
205709 06/17/07 03:28 A. Camsonne Radiator temperature
205708 06/17/07 02:41 A. Camsonne Radiator out
205705 06/17/07 02:18 A. Camsonne Raster run 1297
205704 06/17/07 02:15 A. Camsonne IOCHA5 down
205698 06/17/07 00:00 X. Zheng swing shift summary
Saturday
205695 06/16/07 23:17 A. Camsonne Dead time looks fine now
205694 06/16/07 22:48 X. Zheng radiator temperature stripchart
205693 06/16/07 22:46 X. Zheng radiation monitor and ozone level
205692 06/16/07 22:42 X. Zheng HV (Right HRS) screen shot
205691 06/16/07 22:41 X. Zheng tools screen
205690 06/16/07 22:37 X. Zheng FPP threshold turned on
205688 06/16/07 22:26 K. Allada Target moved to 3He(loop1)
205687 06/16/07 22:19 X. Zheng radiator temperature slope
205686 06/16/07 22:16 X. Zheng radiator in and target to 3He
205681 06/16/07 21:58 K. Allada Optics plots for run 1291
205679 06/16/07 21:36 X. Zheng radiator was out since begining of this shift
205677 06/16/07 21:28 K. Allada Target moved to carbon
205675 06/16/07 21:26 X. Zheng raster setting: 3.0 by 3.9 (MCC)
205673 06/16/07 21:22 X. Zheng Hole!
205670 06/16/07 20:05 X. Zheng harp scan result (plot)
205662 06/16/07 19:27 X. Zheng high deadtime
205658 06/16/07 19:17 X. Zheng left detector HRS IOC fixed
205656 06/16/07 19:03 K. Allada Left arm TS11 rebooted
205654 06/16/07 18:55 atrig Trigger-Download
205648 06/16/07 18:29 X. Zheng access postponed
205647 06/16/07 18:28 K.Allada Target moved to carbon hole
205646 06/16/07 18:24 X. Zheng Harp scan results
205645 06/16/07 17:22 X. Zheng controlled access
205644 06/16/07 16:44 X. Zheng 50uA in the Hall
205642 06/16/07 16:09 eschulte Day Shift Log
205641 06/16/07 16:03 eschulte iocha4 down, can't access on-line reboot
205640 06/16/07 15:55 eschulte Follow-Up Q3 Message
205639 06/16/07 15:54 eschulte HRS-L Q3 Cycled
205638 06/16/07 15:54 xqian target status
205637 06/16/07 15:53 eschulte MCC: Sending low current CW
205636 06/16/07 15:29 eschulte HRS-L Q3 keeps drifting low
205635 06/16/07 15:15 eschulte MCC: Sending Tune Beam
205634 06/16/07 13:38 eschulte Hall A Tools after Spectrometer Change
205633 06/16/07 13:37 eschulte Day Shift Checklist: Hall A Tools Screen
205632 06/16/07 13:20 A. Camsonne
205631 06/16/07 13:13 eschulte MCC to bring Hall back to Beam Permit
205630 06/16/07 12:51 eschulte Day Shift Checklist: Radiator Temperature
205629 06/16/07 12:51 eschulte Day Shift Checklist: Radiation Monitors
205628 06/16/07 12:50 eschulte Day Shift Checklist: Hall Ozone Monitor
205627 06/16/07 12:49 eschulte Day Shift Checklist: HRS-L HV Screens
205626 06/16/07 12:48 eschulte Day Shift Checklist: HRS-R HV Screens
205625 06/16/07 12:47 sawatzky Right S2 left PMT comment
205624 06/16/07 12:19 A. Camsonne Run 1279 pedestal run
205621 06/16/07 12:06 A. Camsonne Right S2 Left paddle 2
205620 06/16/07 12:00 A. Camsonne OR of S1 and OR of S2 removed from Gas Cerenkov/S0 discriminator
205619 06/16/07 11:58 atrig Trigger-Download
205616 06/16/07 11:26 xqian target IOC reboot
205607 06/16/07 10:52 atrig Trigger-Download
205595 06/16/07 09:15 eschulte ARM: Survey Results
205593 06/16/07 08:46 eschulte ARM: Hall in controlled access
205589 06/16/07 08:42 A. Camsonne pedestal subtraction off for checks
205588 06/16/07 08:41 eschulte HRS-L: Q3 drift
205587 06/16/07 08:34 eschulte MCC: Pass change
205585 06/16/07 08:16 eschulte MCC Call: Setup plan for pass change
205583 06/16/07 08:00 V. Sulkosky Owl Shift Summary
205580 06/16/07 06:53 khem high voltage screens
205579 06/16/07 06:52 khem ozone & radiation stripchart
205578 06/16/07 06:44 V. Sulkosky Radiator IN
205575 06/16/07 05:38 V. Sulkosky Radiator OUT
205570 06/16/07 04:25 A. Camsonne Target screens (2)
205569 06/16/07 04:14 V. Sulkosky Radiator Temperature Stripchart
205566 06/16/07 03:18 A. Camsonne Target screens
205565 06/16/07 02:25 V. Sulkosky Hall A Tools and BCM Chart
205562 06/16/07 01:46 A. Camsonne Right S2 left PMT
205561 06/16/07 01:12 V. Sulkosky Radiator IN
205558 06/16/07 00:06 V. Sulkosky Raster Check 1257
Friday
205555 06/15/07 23:50 A. Camsonne Moving from C to He3 target
205553 06/15/07 23:45 emily Swing Shift Summary
205550 06/15/07 22:26 jglister Ozone stripchart
205549 06/15/07 22:25 jglister Radiator temp stripchart
205548 06/15/07 22:24 jglister Hall A Tools Screen
205547 06/15/07 22:24 jglister High voltage screens 2
205546 06/15/07 22:23 jglister High voltage screens 1
205541 06/15/07 20:17 jglister Target screens
205540 06/15/07 19:55 emily replay of 1251
205535 06/15/07 18:19 jglister Target to carbon
205528 06/15/07 15:58 Saha Shift Summary
205526 06/15/07 14:16 Xiaohui Target Screen: 3He
205525 06/15/07 14:15 Xiaohui Target Screen
205518 06/15/07 11:24 Yi Qiang Raster check
205514 06/15/07 09:26 Xiaohui Beam Stopped
205509 06/15/07 08:04 V. Sulkosky Owl Shift Summary
205508 06/15/07 07:58 V. Sulkosky Radiator IN
205505 06/15/07 07:11 V. Sulkosky Radiator OUT
205502 06/15/07 06:24 V. Sulkosky Right Arm HV Screens
205501 06/15/07 06:23 V. Sulkosky Left Arm and Beamline HVs
205498 06/15/07 05:28 V. Sulkosky Radiator Temperature
205497 06/15/07 05:22 V. Sulkosky Beamline HV Alarm
205494 06/15/07 04:33 Chiranjib Target screens II
205493 06/15/07 04:32 Chiranjib Target screens I
205492 06/15/07 04:14 V. Sulkosky Radiation and Ozone Stripcharts
205489 06/15/07 02:47 V. Sulkosky Beam is Back
205488 06/15/07 02:46 V. Sulkosky Raster Check 1237
205486 06/15/07 01:55 V. Sulkosky No Beam
205485 06/15/07 01:53 V. Sulkosky Radiator IN
205484 06/15/07 01:51 Chiranjib Target----> He-3
205482 06/15/07 00:47 V. Sulkosky Hall A Tools and BCM Chart
205479 06/15/07 00:12 Chiranjib Target screens II
205478 06/15/07 00:11 Chiranjib Target Screens I
Thursday
205476 06/14/07 23:54 emily Swing Shift Summary
205471 06/14/07 22:10 jglister High voltage screens 2
205470 06/14/07 22:09 jglister High voltage screens 1
205469 06/14/07 22:09 jglister Ozone stripchart
205468 06/14/07 22:09 jglister Radiator Temp stripchart
205467 06/14/07 22:08 jglister Hall A Tools screen
205462 06/14/07 20:00 jglister Target screens
205454 06/14/07 17:12 jglister Target to carbon
205451 06/14/07 16:00 Yi Qiang Day shift summary
205450 06/14/07 15:54 Xiaohui Spot++ run 1227
205448 06/14/07 15:52 Yi Qiang Radiator IN and beam is back !!
205446 06/14/07 15:36 Yi Qiang Beam down
205443 06/14/07 14:52 Yi Qiang Radiator temperature
205442 06/14/07 14:52 Yi Qiang Hall A tools
205440 06/14/07 14:32 Yi Qiang Radiator OUT
205432 06/14/07 10:16 Yi Qiang Radiator IN
205428 06/14/07 09:04 Yi Qiang Radiator OUT
205425 06/14/07 08:36 Yi Qiang Beam down for about 10 minutes
205420 06/14/07 08:03 V. Sulkosky Owl Shift Summary
205417 06/14/07 07:17 V. Sulkosky Radiator Temperature Chart
205416 06/14/07 06:08 V. Sulkosky Right Arm HV
205415 06/14/07 06:07 V. Sulkosky Left Arm and Beamline HV
205412 06/14/07 05:05 V. Sulkosky T4 Prescale Factor Adjusted
205411 06/14/07 04:54 V. Sulkosky Radiator IN
205408 06/14/07 04:25 V. Sulkosky Radiation Stripchart
205407 06/14/07 03:43 V. Sulkosky Raster Check 1214
205406 06/14/07 03:40 V. Sulkosky Radiator OUT
205403 06/14/07 02:54 V. Sulkosky Spot++ Warnings II
205400 06/14/07 02:09 V. Sulkosky Hall A Tools and BCM Chart
205395 06/14/07 01:02 V. Sulkosky Ozone StripChart
205394 06/14/07 00:43 V. Sulkosky Spot++ Warnings
205393 06/14/07 00:42 V. Sulkosky Raster Check 1210
205390 06/14/07 00:06 Xinhu Yan Target screen
Wednesday
205388 06/13/07 23:59 emily Swing Shift Summary
205383 06/13/07 21:39 jglister Radiator Temp stripchart
205382 06/13/07 21:39 jglister Ozone stripchart
205381 06/13/07 21:38 jglister Hall A Tools screen
205380 06/13/07 21:38 jglister High voltage screens 2
205379 06/13/07 21:37 jglister High voltage screens 1
205376 06/13/07 20:19 jglister Target screens
205374 06/13/07 20:09 jglister Target to He3
205372 06/13/07 19:57 emily Spot++
205370 06/13/07 19:33 jglister Target to carbon hole
205368 06/13/07 16:31 sarty HV for HRS-L Pion Rejector 2: turned off
205364 06/13/07 16:02 R. Michaels day shift summary
205363 06/13/07 14:00 R. Michaels carbon run 1203
205357 06/13/07 10:11 A. Camsonne S0 peaks and crosstalk
205356 06/13/07 09:42 R. Michaels S0 multiple TDC peaks
205355 06/13/07 09:09 Adaq radiator out
205354 06/13/07 09:03 R. Michaels target moved to carbon.
205352 06/13/07 08:04 V. Sulkosky Owl Shift Summary
205348 06/13/07 07:23 V. Sulkosky Right arm HV Screens
205347 06/13/07 07:22 V. Sulkosky Left Arm and Beamline HV
205346 06/13/07 07:20 V. Sulkosky Radiation Stripchart
205345 06/13/07 06:57 V. Sulkosky Radiator IN
205342 06/13/07 05:46 V. Sulkosky Radiator OUT
205337 06/13/07 04:44 V. Sulkosky TECH: Ozone Stripchart
205336 06/13/07 04:30 V. Sulkosky Beam back!
205334 06/13/07 04:07 V. Sulkosky No Beam
205329 06/13/07 03:08 V. Sulkosky Radiator Temperature
205326 06/13/07 02:12 V. Sulkosky Run 1192 End-of-Run
205322 06/13/07 01:42 V. Sulkosky Hall A Tools and BCM Chart
205321 06/13/07 00:30 V. Sulkosky Beam Back!
205319 06/13/07 00:26 V. Sulkosky Beam Off
205317 06/13/07 00:15 V. Sulkosky Raster for Run 1191
205316 06/13/07 00:07 Xinhu Yan Target snapshot
Tuesday
205314 06/12/07 23:57 emily Swing Shift Summary
205312 06/12/07 23:32 jglister Target to He3
205306 06/12/07 22:17 jglister Ozone stripchart
205305 06/12/07 22:16 jglister Hall A Tools screen
205304 06/12/07 22:16 jglister Radiator Temp stripchart
205303 06/12/07 22:15 jglister High voltage screens 2
205302 06/12/07 22:14 jglister High voltage screens 1
205292 06/12/07 20:36 jglister Target to Dummy
205290 06/12/07 20:06 jglister Target screens
205278 06/12/07 15:59 Yi Qiang Day shift summary
205277 06/12/07 15:12 Xiaohui Run 1177: Radiator out
205275 06/12/07 15:09 Yi Qiang Radiator IN
205273 06/12/07 14:37 xqian Target Status
205271 06/12/07 13:54 Yi Qiang Radiator OUT
205269 06/12/07 13:16 Yi Qiang Radiation and Ozone monitors
205268 06/12/07 13:14 Yi Qiang Radiator Temperature
205267 06/12/07 13:14 Yi Qiang HV screen
205266 06/12/07 13:13 Yi Qiang Hall A Tools Screen
205261 06/12/07 12:42 Xiaohui Spot++ run 1174
205255 06/12/07 12:24 A. Camsonne S0 check
205254 06/12/07 10:54 sawatzky Runs 1169--1171 are with 6% radiator
205251 06/12/07 10:18 sawatzky Radiator IN (starting with run 1169)
205243 06/12/07 08:01 V. Sulkosky Owl Shift Summary
205242 06/12/07 07:26 Xinhu Yan Target snapshot
205241 06/12/07 07:18 V. Sulkosky Radiator OUT
205240 06/12/07 07:05 V. Sulkosky Raster check 1168
205234 06/12/07 05:48 V. Sulkosky HV Screens II
205233 06/12/07 05:48 V. Sulkosky HV Screens I
205232 06/12/07 05:45 V. Sulkosky Radiation Stripchart
205231 06/12/07 04:41 V. Sulkosky Beam Back
205228 06/12/07 04:19 V. Sulkosky Beam Off
205227 06/12/07 03:39 V. Sulkosky BCM and Ozone Charts
205224 06/12/07 02:52 V. Sulkosky Radiator Temperature
205221 06/12/07 01:50 V. Sulkosky Raster Size
205218 06/12/07 01:21 V. Sulkosky Late Entry: Rad. IN
205215 06/12/07 00:39 V. Sulkosky BPM Position Charts
205214 06/12/07 00:29 V. Sulkosky Beam position
205213 06/12/07 00:18 V. Sulkosky Hall A Tools
205212 06/12/07 00:11 Xinhu Yan Target Screen
Monday
205210 06/11/07 23:46 sarty Swing Shift Summary
205203 06/11/07 20:55 jglister High voltage screens 2
205202 06/11/07 20:54 jglister High voltage screens 1
205201 06/11/07 20:53 jglister Ozone stripchart
205200 06/11/07 20:53 jglister Radiator temp stripchart
205199 06/11/07 20:52 jglister Hall A Tools screen
205196 06/11/07 20:09 sarty Radiator temp stable @ ~ 110 deg C
205195 06/11/07 19:54 jglister Target screens
205190 06/11/07 17:50 sarty Initial radiator heating (20 uA, 3.2 GeV)
205186 06/11/07 17:39 jglister Target to He3
205185 06/11/07 17:36 emily Radiator IN
205184 06/11/07 17:36 sarty Beam size @ 3.2 GeV
205181 06/11/07 17:15 jglister Target to optics
205178 06/11/07 17:07 jglister Target to BeO
205177 06/11/07 17:06 Adaq spot++
205172 06/11/07 16:55 Adaq Spot++ on carbon hole tgt
205170 06/11/07 16:47 sawatzky Harp scan results
205166 06/11/07 16:27 jglister Target to carbon hole
205164 06/11/07 16:02 Yi Qiang Day shift summary
205163 06/11/07 15:52 Yi Qiang Kinematics Setting Changed
205162 06/11/07 15:49 xqian Target status
205160 06/11/07 15:23 Xiaohui FPP chamber 3 wire hits
205158 06/11/07 14:27 Yi Qiang Move radiator OUT
205157 06/11/07 14:25 gilman long time scale view of rad temperature
205155 06/11/07 13:23 Yi Qiang Ozone chart updated
205152 06/11/07 12:47 xqian HALLA tools
205151 06/11/07 12:47 xqian ozone stripchart
205150 06/11/07 12:46 xqian radiator temperature snapshot
205149 06/11/07 12:46 xqian HV snapshot
205148 06/11/07 12:20 Xiaohui FPP wire group replay
205147 06/11/07 12:14 Jack Gas supply problem this morning
205145 06/11/07 12:02 Xiaohui Spot++ 1143
205143 06/11/07 12:00 Xiaohui run 1142: cosmic
205140 06/11/07 11:41 M. Stevens, Ozone monitor restored to operation
205139 06/11/07 10:36 Xiaohui VDC efficiency replay
205137 06/11/07 09:03 Yi Qiang Beam down
205133 06/11/07 08:54 Yi Qiang Argon bottle changed
205132 06/11/07 08:06 A. Camsonne Owl shift summary
205129 06/11/07 07:46 A. Camsonne Hall C is up beam back in Hall A
205128 06/11/07 07:46 A. Camsonne Radiator temperature
205127 06/11/07 07:29 A. Camsonne Radiation monitors
205126 06/11/07 07:29 A. Camsonne Ozone
205125 06/11/07 07:28 A. Camsonne HV Right arm
205124 06/11/07 07:28 A. Camsonne HV left arm
205123 06/11/07 07:26 A. Camsonne Hall A tools
205122 06/11/07 07:18 A. Camsonne Beam off to troubleshoot Hall C
205115 06/11/07 06:37 A. Camsonne Beam back going to 30 uA to put radiator in
205114 06/11/07 06:26 A. Camsonne MCC working on Hall C Beam off for a bit
205109 06/11/07 05:07 A. Camsonne Beam to Hall A
205108 06/11/07 03:40 A. Camsonne No beam for a few minutes
205105 06/11/07 03:20 byung target -> carbon
205103 06/11/07 03:07 A. Camsonne Radiator out
205100 06/11/07 02:31 A. Camsonne Cycled Q3 during target change
205099 06/11/07 02:29 byung target chart2
205098 06/11/07 02:28 byung target chart
205097 06/11/07 02:28 byung target -> dummy
205088 06/11/07 00:06 Sarty Swing Shift Summary
Sunday
205084 06/10/07 22:52 jglister Ozone Stripchart
205083 06/10/07 22:50 jglister Radiator Temp Stripchart
205082 06/10/07 22:49 jglister High Voltage Screens 2
205081 06/10/07 22:48 jglister High Voltage Screens 1
205080 06/10/07 22:46 jglister Hall A Tools screen
205079 06/10/07 22:11 sarty Radiator IN: temp seems ok with 30 uA
205067 06/10/07 20:25 jglister Target to He3
205062 06/10/07 19:50 jglister Target to optics
205061 06/10/07 19:45 Sarty Raster set for 2.1 GeV (MCC "3.5 x 4.5" gives 2x2)
205056 06/10/07 19:15 Sarty Beam spot size ok @ 2.1 GeV
205055 06/10/07 18:48 jglister Target screens
205050 06/10/07 15:53 X. Zheng day shift summary
205049 06/10/07 15:35 sawatzky Installed 1.5+9" FPP doors (now: 1.5+3+6+9)
205048 06/10/07 15:05 X. Zheng radiator out
205047 06/10/07 15:02 X. Zheng radiator temperature stripchart
205046 06/10/07 15:02 X. Zheng radition monitor
205045 06/10/07 15:01 X. Zheng Tools screen
205044 06/10/07 13:35 X. Zheng pulsed beam
205043 06/10/07 12:24 Adaq target was changed to carbon hole
205042 06/10/07 12:15 A. Camsonne Retiming signal check for T2
205041 06/10/07 11:59 X. Zheng radiator temperature going up (no beam)
205040 06/10/07 11:53 sawatzky Radiator temp "rebound" without beam
205039 06/10/07 11:37 X. Zheng beam status update
205038 06/10/07 11:14 atrig Trigger-Download
205037 06/10/07 11:10 X. Zheng back to beam permit
205036 06/10/07 10:40 X. Zheng Both HRS momentum settled
205035 06/10/07 10:28 atrig Trigger-Download
205034 06/10/07 10:28 X. Zheng controlled access
205033 06/10/07 09:49 X. Zheng Left HRS momentum settled at the new setting
205031 06/10/07 09:35 X. Zheng Spectrometer angle changed to 53.690 degrees
205030 06/10/07 09:32 sawatzky BPM stripchart
205029 06/10/07 09:24 sawatzky Radiation and Ozone stripchart
205028 06/10/07 09:19 sawatzky Radiator temp plot for 30uA running
205023 06/10/07 07:59 gron Owl Shift Summary
205020 06/10/07 07:04 Adaq shift summery
205018 06/10/07 07:01 adaq ROC 1 reboot
205015 06/10/07 06:37 adaq spot++ raster
205011 06/10/07 06:18 gron Moved Spectrometers to 63.62 degrees
204999 06/10/07 05:11 Adaq raster spot++
Saturday
204977 06/09/07 23:56 Ramesh Shif Summary
204974 06/09/07 23:18 sawatzky Raster scan "results"
204973 06/09/07 23:16 sawatzky Radiator temp for 30 uA running
204972 06/09/07 23:09 sawatzky BPM stripcharts
204969 06/09/07 23:02 sawatzky Raster scan: 12x18 (MCC), 8x8mm
204966 06/09/07 22:48 sawatzky Raster scan: 6x8 (MCC units), 5x6 mm
204964 06/09/07 22:40 sawatzky Raster scan: 3x4 (MCC units)
204962 06/09/07 22:36 byung target chart2
204961 06/09/07 22:36 byung target chart
204957 06/09/07 22:00 sawatzky tech: ozone montor ioc is dead (iocha12?)
204954 06/09/07 21:56 sawatzky ROC2 fixed after controlled access
204953 06/09/07 21:27 sawatzky Radiator Temps @ 10uA and 30uA
204952 06/09/07 21:27 sawatzky ROC2 died
204949 06/09/07 21:07 sawatzky spot++ for Run 1086
204946 06/09/07 20:35 byung target chart2
204945 06/09/07 20:35 byung target chart
204936 06/09/07 20:22 sawatzky Radiator IN (6%)
204935 06/09/07 20:22 byung target -> loop1
204934 06/09/07 20:17 Ramesh Holey target
204932 06/09/07 20:02 sawatzky Target to Carbon Hole
204928 06/09/07 19:48 sawatzky Target to BeO
204919 06/09/07 19:24 sawatzky Where's the hole?
204915 06/09/07 19:11 sawatzky Harp scan results: 124x111 um
204911 06/09/07 18:49 sawatzky Radiator out for beam setup
204907 06/09/07 18:41 byung target change -> 'C hole'
204906 06/09/07 18:36 sawatzky Harp scan
204902 06/09/07 17:47 E.Chudakov Moller results
204896 06/09/07 14:37 E.Chudakov MCC started tuning for Moller
204895 06/09/07 13:44 E.Chudakov Wien angle
204893 06/09/07 12:30 Jack Left Arm Shield House doors - right door is disabled
204892 06/09/07 12:24 Jack VME crate for iocha14 replaced
204883 06/09/07 10:07 atrig Trigger-Download
204882 06/09/07 10:07 A. Camsonne Setting T4 or of S1 or S2
204881 06/09/07 09:28 adaq ROC 3 reboot
204875 06/09/07 08:48 adaq tech
204874 06/09/07 08:43 gilman shift summary
204871 06/09/07 08:03 A. Camsonne iocha14 down
204869 06/09/07 07:26 A. Camsonne S0 HV
204866 06/09/07 07:17 A. Camsonne S0A HV increased to 1850V
204862 06/09/07 07:08 eip/ab radiator temperature 20 micro amp
204858 06/09/07 06:26 A. Camsonne Coincidence peak run 1053
204850 06/09/07 03:56 A. Camsonne Target hole
204844 06/09/07 03:37 gilman FPP status
204835 06/09/07 03:12 atrig Trigger-Download
204832 06/09/07 02:56 A. Camsonne Carbon hole run 1043
Friday
204806 06/08/07 23:53 Ramesh Shift summary
204804 06/08/07 23:26 atrig Trigger-Download
204803 06/08/07 23:24 atrig Trigger-Download
204800 06/08/07 23:01 atrig Trigger-Download
204799 06/08/07 22:39 atrig Trigger-Download
204797 06/08/07 22:20 A. Camsonne Pedestal subtraction back on
204796 06/08/07 22:19 A. Camsonne S0 T2 trigger
204794 06/08/07 22:07 sawatzky Work around for LHRS hut doors
204793 06/08/07 21:48 atrig Trigger-Download
204790 06/08/07 21:24 Ramesh Target charts
204789 06/08/07 21:20 Ramesh Detector high voltages
204788 06/08/07 21:16 Ramesh Hall A tools
204787 06/08/07 21:12 atrig Trigger-Download
204786 06/08/07 21:08 atrig Trigger-Download
204785 06/08/07 21:03 sawatzky Left shield hut doors not working again
204783 06/08/07 20:54 A. Camsonne Pedestal subtraction off for run 1027
204778 06/08/07 19:42 Jack Left Shield House doors
204774 06/08/07 19:28 Jack Carbon fiber covers for the Right Arm VDCs installed
204773 06/08/07 19:25 Jack Right Arm top VDC threshold
204767 06/08/07 17:27 folts tech
204765 06/08/07 17:19 sawatzky RHRS floor angle marks corrected
204764 06/08/07 17:05 atrig Trigger-Download
204763 06/08/07 16:32 sawatzky/gil Radiator stepper motor voltage settings
204752 06/08/07 15:24 ewing checklist
204745 06/08/07 14:58 A. Beck Quades cycling
204740 06/08/07 13:45 gilman HRS right floor marks
204739 06/08/07 13:39 A. Camsonne/ S0 right arm cabling
204714 06/08/07 11:02 A. Camsonne E04018 backup
204713 06/08/07 10:46 A. Camsonne Ozone
204712 06/08/07 10:43 A. Camsonne Radiation monitors
204710 06/08/07 10:04 A. Camsonne Day shift summary
204704 06/08/07 08:03 B.Hahn Owl Shift Summary
204701 06/08/07 07:34 B.Hahn Spot++ for run 4182
204699 06/08/07 06:28 kelleher target screens
204698 06/08/07 06:27 B.Hahn Spot++ for run 4181
204696 06/08/07 06:12 kelleher target to dummy
204693 06/08/07 05:35 B.Hahn False alarm
204692 06/08/07 05:30 B.Hahn Fire Alarm
204690 06/08/07 05:11 B.Hahn Spot++ for run 4179
204687 06/08/07 04:51 B.Hahn Spot++ for run 4178
204682 06/08/07 04:05 B.Hahn Spot++ for run 4176
204677 06/08/07 03:25 B.Hahn RHRS HV Screen
204676 06/08/07 03:24 B.Hahn LHRS HV Screen
204675 06/08/07 03:23 B.Hahn Ozone and Radiation Monitor
204674 06/08/07 03:21 B.Hahn Hall A Tools and Beam Current Monitor
204668 06/08/07 02:21 B.Hahn Spot++ for run 4172
204665 06/08/07 02:08 B.Hahn spot++ for run 4116
204663 06/08/07 01:43 atrig Trigger-Download
204628 06/08/07 00:06 R.Lindgren Swing Shift Summary
Thursday
204623 06/07/07 22:51 Lindgren Hall A Tools
204622 06/07/07 22:50 R.Lindgren Beam Current Plot
204610 06/07/07 22:14 V. Sulkosky Collimators IN Runs 4144 and 4145
204604 06/07/07 21:57 atrig Trigger-Download
204603 06/07/07 21:48 A. Camsonne Helicity signal
204595 06/07/07 19:40 Jack Buzz coming from the Detectors Gas System Interlock Status Box
204584 06/07/07 17:34 Lindgren Starting Swing Shift
204583 06/07/07 17:33 Lindgren Hall A Tools
204576 06/07/07 16:34 Xinhu Yan Target Screen beam 90 uA
204575 06/07/07 16:21 Higinbotham Shift Summary
204571 06/07/07 15:57 Higinbotham Current Screen
204563 06/07/07 14:17 Ida Rodrigue Raster
204562 06/07/07 14:07 Piotr Decows PS3 changed from 2 to 1
204559 06/07/07 13:56 Ida Rodrigue Hall A Tools
204556 06/07/07 13:03 Piotr Decows Target charts at I=100uA
204554 06/07/07 12:19 jpchen target operating temp changed to be 8.6K
204553 06/07/07 12:09 Elena Kh. Current monitor
204552 06/07/07 11:49 jpchen target charts
204551 06/07/07 11:48 jpchen target charts at 8.0 K, 200 psi
204550 06/07/07 11:44 Piotr Decows Control access
204547 06/07/07 11:16 Piotr Decows Moving target
204539 06/07/07 08:48 Ida Rodrigue Target Controls
204538 06/07/07 08:46 Piotr Decows Target charts
204535 06/07/07 08:01 B.Hahn Owl Shift Summary
204533 06/07/07 07:34 B.Hahn spot++ for run 4118
204528 06/07/07 05:54 kelleher target screens
204527 06/07/07 05:53 B.Hahn spot++ for run 4116
204522 06/07/07 04:24 B.Hahn TECH Alarm coming from "Detectors Gas System Interlock Status"
204521 06/07/07 04:14 B.Hahn Tools Screen
204518 06/07/07 03:31 B.Hahn Spot++ for Run 4113
204517 06/07/07 03:28 B.Hahn Beamline HV alarm
204514 06/07/07 03:04 B.Hahn RHRS HV Screens
204513 06/07/07 03:04 B.Hahn LHRS HV Screen
204512 06/07/07 03:02 B.Hahn Ozone and Radiation Monitor
204511 06/07/07 03:01 B.Hahn Hall A Tools and Beam Current Monitor
204510 06/07/07 03:00 B.Hahn Spot++ for run 4112
204507 06/07/07 02:04 B.Hahn spot++ for run 4111
204502 06/07/07 00:32 B.Hahn spot++ for run 4109
204499 06/07/07 00:09 B.Hahn spot++ for run 4108
204498 06/07/07 00:01 Sparveris Shift summary
Wednesday
204496 06/06/07 23:41 Sparveris beam is back
204494 06/06/07 23:34 Sparveris RF problem
204490 06/06/07 22:19 Xinhu Yan Target alarm
204489 06/06/07 22:09 Xinhu Yan Target alarm
204480 06/06/07 18:38 Sparveris Tools screen
204479 06/06/07 18:37 Sparveris beam is back on
204477 06/06/07 17:24 Sparveris no beam
204473 06/06/07 16:27 Sparveris spot++
204472 06/06/07 16:14 Xinhu Yan Target Charts
204470 06/06/07 16:08 R. Lindgren Shift Summary
204469 06/06/07 15:57 R. Michaels except, plan to install S0 on R-arm
204466 06/06/07 15:50 R. Michaels Need no trigger timing mods for e03101
204465 06/06/07 15:39 gilman re-adjusted density
204464 06/06/07 15:27 gilman adjust cell properties
204461 06/06/07 14:31 gilman cryotarget pid settnig changed
204460 06/06/07 14:28 gilman changed cell parameters
204457 06/06/07 14:10 Lindgren Hall A Tools with Beam
204456 06/06/07 14:08 gilman target snapshot
204455 06/06/07 14:08 Lindgren spot++
204452 06/06/07 13:10 Lindgren Hall A Tools
204451 06/06/07 12:56 Lindgren Beam Back
204449 06/06/07 12:12 Lindgren Beam Off
204447 06/06/07 11:55 Lindgren Hall A Tools
204443 06/06/07 11:10 Lindgren Beam Loss Monitor
204442 06/06/07 10:45 R. Michaels change scaler VME software for helicity in-time
204439 06/06/07 09:58 R. Lindgren FSD card
204438 06/06/07 08:27 R. Lindgren Beam not available
204437 06/06/07 08:04 Piotr Decows Re: target fluctuations
204436 06/06/07 08:03 Elena Kh. Shift Summary
204434 06/06/07 06:43 Elena Kh. beam off
204430 06/06/07 05:37 Elena Kh. RHRS HV Screens
204429 06/06/07 05:36 Elena Kh. LHRS HV Screen
204428 06/06/07 05:35 Elena Kh. Radiational and Ozone monitors
204427 06/06/07 05:34 Elena Kh. Current Monitor and Target BPM chart
204426 06/06/07 05:33 Elena Kh. Hall A General Tools
204425 06/06/07 05:25 Elena Kh. spot++ for run 4091
204419 06/06/07 04:13 Elena Kh. spot++ for run 4089
204414 06/06/07 02:23 Elena Kh. spot++ for run 4087
204413 06/06/07 02:10 Piotr Decows Change of PS1
204412 06/06/07 02:02 Piotr Decows Target fluctuations, cont.
204409 06/06/07 01:51 Piotr Decows Target fluctuations, cont.
204408 06/06/07 01:49 Elena Kh. Radiation and Ozone monitors
204405 06/06/07 01:08 Elena Kh. General Accelerator Screen
204404 06/06/07 01:06 Elena Kh. Hall A General Tools
204403 06/06/07 01:06 Elena Kh. RHRS HV Screen
204402 06/06/07 01:05 Elena Kh. LHRS HV Screen
204401 06/06/07 01:04 Elena Kh. Hall A Current Monitor and TargetBPM chart
204400 06/06/07 01:01 Elena Kh. spot++ for run 4085
204399 06/06/07 00:58 Piotr Decows Target controls
204398 06/06/07 00:56 Piotr Decows Target temp/pressure fluctuations
204394 06/06/07 00:07 bogdanw shift summary
Tuesday
204391 06/05/07 22:30 Xinhu Yan Target Moved to 4He
204386 06/05/07 22:07 bogdanw target position check
204383 06/05/07 21:58 bogdanw raster for run 4080
204371 06/05/07 19:07 bogdanw HC HRS-right
204370 06/05/07 19:03 bogdanw HRSs for He4 77 fm-2
204369 06/05/07 19:03 Jack Work during the access
204367 06/05/07 18:32 jpchen Target operation T increased to 8.1K.
204366 06/05/07 17:39 jpchen Target charts at 7.5K, 200 psi
204363 06/05/07 16:45 Xinhu Yan Target Screen
204361 06/05/07 16:23 atrig Trigger-Download
204360 06/05/07 16:14 gilman day shift summary
204357 06/05/07 15:53 Piotr Decows Target moved to empty
204356 06/05/07 15:20 gilman cameras, q2, q3 left
204355 06/05/07 15:19 atrig Trigger-Download
204354 06/05/07 15:18 atrig Trigger-Download
204353 06/05/07 14:57 gilman Hall A tools screen
204349 06/05/07 12:50 Piotr Decows Target controls
204348 06/05/07 12:48 Piotr Decows Target charts
204343 06/05/07 11:15 A. Saha/A. C BCM U3 got unplugged, replugged now
204336 06/05/07 10:05 R. Michaels more on BCM discrepancies
204335 06/05/07 09:41 Piotr Decows Target controls
204334 06/05/07 09:40 Piotr Decows Target charts
204333 06/05/07 09:39 R. Michaels re: BCM discrepancies
204332 06/05/07 09:30 A. Camsonne BCM u3
204328 06/05/07 09:15 A. Camsonne BCM u10
204327 06/05/07 09:14 gilman beam coming back
204326 06/05/07 08:38 B. Hahn Beam Down a while longer
204325 06/05/07 08:01 M.Olson Owl Shift Summary
204323 06/05/07 07:18 B. Hahn Beam down a while
204322 06/05/07 07:12 B. Hahn Beam down
204321 06/05/07 06:58 B. Hahn Target Screens II
204320 06/05/07 06:57 B. Hahn Target Screens I
204316 06/05/07 05:28 M.Olson spot++ for run 4061
204311 06/05/07 04:28 M.Olson spot++ for run 4059
204310 06/05/07 03:51 M.Olson spot++ for run 4058
204307 06/05/07 03:25 M.Olson RHRS HV Screens
204306 06/05/07 03:24 M.Olson LHRS HV Screen
204305 06/05/07 03:24 M.Olson Ozone and Radiation Monitor
204304 06/05/07 03:21 M.Olson Hall A Tools and Beam Current Monitor
204303 06/05/07 03:10 M.Olson Discrepencies in BCM values in endrun summaries
204301 06/05/07 03:00 M.Olson Beam back
204300 06/05/07 02:38 M.Olson Update on beam down
204298 06/05/07 02:29 M.Olson MCC reports problem with capture system (injector) 10-15mins
204297 06/05/07 02:29 B. Hahn Target Screens II
204296 06/05/07 02:28 B. Hahn Target Screens I
204291 06/05/07 00:52 M.Olson Beam back on at 65uA (spot++ for run 4055)
204289 06/05/07 00:02 Armando Acha Swing Shift Summary
Monday
204286 06/04/07 23:35 A. Acha ozone level
204285 06/04/07 23:34 A. Acha quite steady beam
204280 06/04/07 22:16 A. Acha HV screens
204279 06/04/07 22:15 A. Acha HV screen
204277 06/04/07 21:43 A. Acha PS3 test
204271 06/04/07 20:14 A. Acha Beam !!
204266 06/04/07 18:31 A. Acha Ozone stripchart
204263 06/04/07 18:07 A. Acha Hall A tools
204262 06/04/07 18:05 A. Acha Beam strip
204257 06/04/07 16:35 Xinhu Yan Target Screen
204256 06/04/07 16:35 Armando Acha target HP heater oscillation
204254 06/04/07 16:07 Lindgren Shift Summary
204253 06/04/07 15:52 Elena Kh. spot++ for run 4044
204252 06/04/07 15:50 Lindgren Beam back
204250 06/04/07 15:34 Lindgren Lost beam
204248 06/04/07 15:10 Elena Kh. spot++ for run 4043
204247 06/04/07 15:05 Lindgren HRS-R
204244 06/04/07 14:39 Elena Kh. Target screens
204243 06/04/07 14:10 Elena Kh. spot++ for run 4042
204238 06/04/07 12:56 Elena Kh. spot++ for run 4040
204235 06/04/07 12:26 Lindgren Midshift summary
204234 06/04/07 12:22 Lindgren Hall A Tools
204233 06/04/07 12:20 Elena Kh. spot++ for run 4039???
204228 06/04/07 10:13 Elena Kh. Target screens
204223 06/04/07 08:19 Lindgren Day Shift Begins
204220 06/04/07 08:03 M.Olson Owl Shift Summary
204218 06/04/07 07:40 M.Olson spot++ for run 4034
204215 06/04/07 06:37 M.Olson Beam back
204214 06/04/07 06:30 M.Olson Hall A Tools and Beam Current Monitor
204213 06/04/07 06:29 M.Olson MCC: Beam down about 15 mins
204212 06/04/07 06:27 M.Olson Radiation and Ozone Monitor
204211 06/04/07 06:26 M.Olson RHRS HV Screens
204210 06/04/07 06:25 M.Olson LHRS HV Screen (this time with the screen)
204209 06/04/07 06:24 M.Olson LHRS HV Screen
204208 06/04/07 06:24 M.Olson spot++ for run 4033
204203 06/04/07 03:59 M.Olson spot++ for run 4031
204198 06/04/07 02:16 M.Olson RHRS HV Screen
204197 06/04/07 02:15 M.Olson LHRS HV Screen
204196 06/04/07 02:15 M.Olson Ozone and Radiation Monitors
204195 06/04/07 02:13 M.Olson Hall A Tools and Beam Current Monitor
204194 06/04/07 02:04 M.Olson spot++ for run 4029
204193 06/04/07 02:02 M.Olson Deadtime should be ~2%, not <10% as stated in run plan (set PS3 = 2)
204188 06/04/07 01:37 M.Olson Note on prescales
204187 06/04/07 01:32 Piotr Decows Target screen
204186 06/04/07 01:29 Piotr Decows Target charts
204185 06/04/07 01:12 M.Olson Incorrect start-run information for 4026 (first run at Q2=68 fm^-2)
204182 06/04/07 00:52 M.Olson spot++ for run 4026 (Q2=68 fm
204180 06/04/07 00:37 atrig Trigger-Download
204178 06/04/07 00:04 Armando Acha Swing Shift Summary
Sunday
204176 06/03/07 23:05 A. Acha spot++
204172 06/03/07 22:38 Armando Acha Target change
204171 06/03/07 22:37 A. Acha beam back
204170 06/03/07 21:01 A. Acha Beam down
204168 06/03/07 20:21 A. Acha Beam away
204165 06/03/07 19:57 A. Acha Hall A tools
204164 06/03/07 19:27 A. Acha beam back
204162 06/03/07 19:01 A. Acha beam down
204161 06/03/07 18:55 Armando Acha Ozone level
204160 06/03/07 18:21 A. Acha No beam yet
204158 06/03/07 18:00 A. Acha beam down
204157 06/03/07 17:42 A. Acha Left Q2
204156 06/03/07 17:28 A. Acha spot check
204155 06/03/07 17:23 Sparveris Target screens
204153 06/03/07 17:03 A. Acha No beam
204152 06/03/07 16:32 V. Sulkosky Max. Current on 4He Target
204151 06/03/07 16:05 bogdanw shift summary
204144 06/03/07 12:28 H.J Lu Beam Current
204143 06/03/07 12:26 H.J Lu HP Heater
204140 06/03/07 11:36 bogdanw HV HRS right
204139 06/03/07 11:36 bogdanw Tools and HV for HRS left
204138 06/03/07 11:35 bogdanw current monitor
204131 06/03/07 08:02 M.Olson Owl Shift Summary
204127 06/03/07 06:17 M.Olson RHRS HV Screens
204126 06/03/07 06:16 M.Olson LHRS HV Screen
204125 06/03/07 06:16 M.Olson Ozone and Radiation Monitors
204124 06/03/07 06:15 M.Olson Hall A Tools and Beam Current Monitor
204123 06/03/07 06:14 Elena Kh. Target screens
204122 06/03/07 05:54 M.Olson spot++ for run 4013
204119 06/03/07 04:48 M.Olson spot++ for run 4012
204114 06/03/07 03:26 M.Olson spot++ for run 4010
204113 06/03/07 03:02 Elena Kh. Target screens
204108 06/03/07 02:13 M.Olson spot++ for run 4008
204107 06/03/07 02:11 M.Olson Hall A Tools and Beam Current Monitor
204106 06/03/07 02:10 M.Olson Ozone and Radiation Monitor
204105 06/03/07 02:08 M.Olson RHRS HV Screens
204104 06/03/07 02:08 M.Olson LHRS HV Screen
204099 06/03/07 01:44 M.Olson Note on max. current for He4 target.
204098 06/03/07 01:14 M.Olson spot++ for run 4006
204097 06/03/07 01:08 M.Olson Asked MCC to recenter beam on BPMs A and B
204091 06/03/07 00:03 Armando Acha Swing Shift Summary
Saturday
204086 06/02/07 22:24 Armando Acha Beam current increase
204083 06/02/07 21:26 Armando Acha ABU
204078 06/02/07 19:16 A. Acha Ozone level
204077 06/02/07 19:12 Armando Acha Beam position
204076 06/02/07 19:06 A. Acha Beam Current
204071 06/02/07 18:35 Sparveris Target screen
204070 06/02/07 18:29 Armando Acha Hall A tools
204069 06/02/07 17:54 A. Acha rebooting ROCs
204068 06/02/07 17:51 A. Acha Run 3995, 3996 and 3997
204067 06/02/07 17:46 A. Acha ROC 2 failure
204064 06/02/07 16:56 A. Acha PS factors for production
204063 06/02/07 16:39 Armando Acha raste check
204062 06/02/07 16:34 Armando Acha Beam position coordinates
204058 06/02/07 16:03 bogdanw shift summary
204055 06/02/07 14:53 H.J Lu Alarm Handler
204052 06/02/07 14:06 H.J Lu Light Off
204051 06/02/07 14:03 H.J Lu light off
204047 06/02/07 13:08 H.J Lu Move Target
204046 06/02/07 12:18 V. Sulkosky No Beam
204045 06/02/07 12:12 V. Sulkosky 4He Target Current Limit
204041 06/02/07 10:36 R. Michaels re: raster on-call
204040 06/02/07 10:33 A. Camsonne Gas cernkov analog sum scaler
204038 06/02/07 10:12 bogdanw HV right arm
204036 06/02/07 10:10 bogdanw ozone and radiation monitors
204035 06/02/07 10:09 bogdanw tools during run 3989
204034 06/02/07 09:48 bogdanw spot++ for run 3989
204027 06/02/07 08:23 bogdanw gas cherenkov rates
204023 06/02/07 07:51 M.Olson Owl Shift Summary
204018 06/02/07 07:19 M.Olson spot++ from run 3983, consistent with runs before power supply swap
204017 06/02/07 06:49 Jack Raster
204015 06/02/07 06:34 M.Olson Raster restored...securing from controlled access
204014 06/02/07 05:35 M.Olson Run Coordinator notified of problem and progress.
204013 06/02/07 05:31 M.Olson Update on raster problem. Jack Segal on his way in to assist.
204012 06/02/07 04:53 M.Olson Lost horizontal raster...(vertical trace on scope)
204010 06/02/07 04:29 M.Olson spot++ for run 3982
204009 06/02/07 04:27 M.Olson Beam recentered on both BPM's
204008 06/02/07 04:21 M.Olson Beam back...asked for recentering in x on BPMA
204006 06/02/07 04:00 M.Olson MCC taking beam for tuning
204004 06/02/07 03:11 M.Olson spot++ for run 3981
204001 06/02/07 02:38 Elena Kh. Target screens
204000 06/02/07 02:36 M.Olson spot++ for run 3980
203997 06/02/07 02:03 M.Olson spot++ for run 3979
203994 06/02/07 01:49 M.Olson RHRS HV Screens
203993 06/02/07 01:47 M.Olson LHRS HV Screen
203992 06/02/07 01:46 M.Olson Ozone and Radiation Monitors
203991 06/02/07 01:43 M.Olson Hall A Tools and Beam Current Monitor
203990 06/02/07 01:39 M.Olson spot++ for run 3978, beam postion (BPM) OK
203986 06/02/07 00:48 jpchen new target operating conditions
203985 06/02/07 00:13 jpchen taregt charts with T=7.5K
203984 06/02/07 00:03 Armando Acha Swing Shift Summary
203982 06/02/07 00:01 jpchen target charts for T=8.5K
Friday
203981 06/01/07 23:52 atrig Trigger-Download
203976 06/01/07 23:05 atrig Trigger-Download
203974 06/01/07 22:49 Armando Acha Harp scan following
203970 06/01/07 22:10 Armando Acha harp scan status
203969 06/01/07 21:12 Armando Acha Beam down
203968 06/01/07 20:37 Armando Acha IMPORTANT
203966 06/01/07 20:33 V. Sulkosky Raster Size
203961 06/01/07 20:24 V. Sulkosky Beam Position (X, Y) = (-4 mm, 1.5 mm)
203959 06/01/07 20:18 V. Sulkosky Beam Position (X,Y) = (-3.5 mm, 1.5 mm)
203956 06/01/07 20:15 V. Sulkosky Beam Position (X,Y) = (-4.5 mm, 1.0 mm)
203952 06/01/07 20:02 V. Sulkosky Raster with Carbon Hole
203947 06/01/07 19:52 Armando Acha Beam position search
203945 06/01/07 19:38 Armando Acha Target position shifting
203938 06/01/07 18:03 Armando Acha MCC report about Beam problems
203936 06/01/07 16:28 Xinhu Yan Target Screen
203935 06/01/07 15:52 Ole Hansen Shift summary
203934 06/01/07 15:39 atrig Trigger-Download
203933 06/01/07 15:37 atrig Trigger-Download
203932 06/01/07 15:32 Piotr Decows BeO target in
203931 06/01/07 15:28 Ole Hansen Run plan: target boiling test
203930 06/01/07 15:14 Piotr Decows Target movement
203929 06/01/07 11:56 Ole Hansen No beam
203921 06/01/07 08:03 M.Olson Owl Shift Summary
203917 06/01/07 07:02 M.Olson spot++ for run 3957
203914 06/01/07 06:05 M.Olson spot++ for run 3956
203911 06/01/07 05:53 R. Michaels target screens
203910 06/01/07 05:52 M.Olson RHRS HV Screens
203909 06/01/07 05:51 M.Olson LHRS HV Screen (this time including the screen)
203908 06/01/07 05:50 M.Olson LHRS HV Screen
203907 06/01/07 05:50 M.Olson Ozone and Radiation Screens
203906 06/01/07 05:49 M.Olson Hall A Tools and Beam Current Monitor
203905 06/01/07 05:10 M.Olson spot++ for run 3955
203902 06/01/07 04:18 M.Olson spot++ for run 3954
203899 06/01/07 03:01 M.Olson spot++ for run 3953
203894 06/01/07 01:51 M.Olson spot++ for run 3951
203893 06/01/07 01:50 M.Olson RHRS HV Screens
203892 06/01/07 01:49 M.Olson LHRS HV Screen
203891 06/01/07 01:48 M.Olson Ozone and Radiation Monitors
203890 06/01/07 01:48 M.Olson Hall A Tools and Beam Current Monitor
203889 06/01/07 01:16 M.Olson Entry 203887 not endrun for 3951, endrun for 3950 not written to halog
203886 06/01/07 01:01 M.Olson spot++ for run 3950
203885 06/01/07 01:00 R. Michaels spot++ had cuts (incorrectly)
203884 06/01/07 00:49 M.Olson Runs 3936-3949 affected by incorrect angle setting on LHRS
203883 06/01/07 00:35 Elena Kh. t3 replay without shower cut
203882 06/01/07 00:04 R. Michaels L-arm angle adjusted (was slightly wrong)
203880 06/01/07 00:03 bogdanw shift summary